OpenStudy (anonymous):

@charlotte123 the sides of a rectangle are increased by a scale factor of 4. smaller rectangle=20cm

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

@Loser66 ! help!

OpenStudy (anonymous):

brb.

OpenStudy (anonymous):

heya

OpenStudy (charlotte123):

Let x and y be the sides of smaller one 2(x + y) = 20 x + y = 10 new sides 4x and 4y new perimeter = 2(4x + 4y) = 8(x + y) = 80 :D or u can just multiply 20 by 4 XD

OpenStudy (anonymous):

so itz 80?

OpenStudy (charlotte123):

yes :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:D

OpenStudy (charlotte123):

HUZZAH ~(*O*)~

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:D

OpenStudy (charlotte123):

anytin else? :D

OpenStudy (anonymous):

yes.

OpenStudy (charlotte123):

link ahead XD